Suppose the short run production function can be represented by Q= 60,000L^2-1000L^3. Then, determine


A) the level labor employement that maximizes the level of output


B) the level of employement that maximizes APL and the maximum APL

1
Expert's answer
2021-12-31T08:54:24-0500

Answer (a).

The total output will be maximized when the marginal product is equal to zero.

"MPL= \frac{dQ}{dL}"

"MPL=120000L-3000L^2=0"

"L(120000-3000L)=0"

"120000-3000L=0"

"3000L=120000"

"L=40"


Answer (b):

The average product is maximized when the marginal product is equal to the average product. Average product is:

"AP=\frac{Q}{L}"

"AP=6000L-1000L^2"

"6000L-1000L^2=120000L-3000L^2"

"4000L^2=60000L"

"4000L=60000"

"L= \frac{60000}{4000}=15"

Then maximum average product is:

"AP=60000\times15-1000\times(15)^2"

"AP=900000-225000"

"AP=675000"